A 3-month Eurodollar futures contract starts with a price of 97.72 and by the time it is sold it is priced at 98.56.  If three contracts are shorted, what is the overall gain or loss?
Answer #1

If the contracts are shorted, then it would be a loss because the price increased.

Loss = (98.56 - 97.72) = 0.84 per share

Loss per contract = 0.84 * 100 = $84

Loss for 3 contracts = 84 * 3 = $252

Can you please upvote? Thank You :-)